The Healthy Favorites Juice Company sells apple juice for 8.3 cents an ounce and raspberry for 9.3 cents and ounce. The company wishes to market and sell 8 ounce cans of apple raspberry juice for 8.7 cents an ounce. How many ounces of each should be mixed?
Answer by ankor@dixie-net.com(22740)

The Healthy Favorites Juice Company sells apple juice for 8.3 cents an ounce and raspberry for 9.3 cents and ounce.
The company wishes to market and sell 8 ounce cans of apple raspberry juice for 8.7 cents an ounce.
How many ounces of each should be mixed?
:
Let x = amt of raspberry required
then result is to be an 8 oz can therefore:
(8-x) = amt of apple required
:
A typical mixture equation
9.3x + 8.3(8-x) = 8.7(8)
9.3x + 66.4 - 8.3x = 69.6
9.3x - 8.3x = 69.6 - 66.4
x = 3.2 oz of raspberry required
then
8 - 3.2 = 4.8 oz of apple
:\
:
Check these values in the original mixture equation